Please help me with this problem, I don't understand how to answer it. Noah says that dividing a number by 2 is the same as multiplying it by 1/2. He wants to know if he is right, and if so, why? How do you respond?
Answer by Theo(13342)

dividing a number by 2 is the same as muliplying a number by 1/2.

that's true.

when you multiply 2 fractions together, you multiply the numerators and then you multiply the denominators and your result is the new fraction.

for example:

1/2 * 3/5 = 1*3 divided by 2 * 5 = 3 / 10

likewise:

also, every number can be expressed as a fraction by dividing it by 1 since any number divided by 1 is equal to the same number.

example:

5 = 5/1

taking the above statements into considerations, if you have 1/2 * x, then it's equivalent to 1/2 * x/1 which you can then multiply together to get 1*x divided by 2*1 which results in x/2.

you get 1/2 * x and x/2 are equivalent.

if you use them on any number, the result will be the same.

example:

8/2 = 4

1/2 * 8 = 8/2 = 4